OpenStudy (miyanna):

x+6/4=3/5 how do I solve?

jagr2713 (jagr2713):

Do you know how to cross multiply?

OpenStudy (miyanna):

Yeah

jagr2713 (jagr2713):

what do you get when you cross multiply it?

OpenStudy (miyanna):

I think x+30/12

jagr2713 (jagr2713):

Ok @juliannescriv....

OpenStudy (miyanna):

Oh yeah now I remember. Thanks

jagr2713 (jagr2713):

After that what do you think you do next?

jagr2713 (jagr2713):

5x+30=12 Then you do what?

jagr2713 (jagr2713):

Seems like you left so 5x+30=12 subtract 30 5x=-18 then you divide x=-18/5